### v2

# Continuum excitations in a spin-supersolid on a triangular lattice

Abstract
Magnetic, thermodynamic, neutron diffraction and inelastic neutron scattering are used to study spin correlations in the easy-axis XXZ triangular lattice magnet K2Co(SeO3)2. Despite the presence of quasi-2D “supersolid” magnetic order, the low-energy excitation spectrum contains no sharp modes and is instead a broad and structured multi-particle continuum. Applying a weak magnetic field drives the system into an $m=1/3$ fractional magnetization plateau phase and restores sharp spin wave modes. To some extent, the behavior at zero field can be understood in terms of spin wave decay. However, the presence of clear excitation minima at the $M$-points of the Brillouin zone suggest that the spinon language may provide a more adequate description, and signals a possible proximity to a Dirac spin liquid state.
The Ising antiferromagnet (AF) on a triangular lattice is the textbook example of geometric frustration [1, 2]. The ubiquitous cartoon shows one spin pointing up, its neighbor pointing down to minimize exchange energy, this configuration leaving the preferred direction for a third spin undefined. Beyond this simplistic picture, the problem is actually a very complex one. The quantum $S=1/2$ nearest-neighbor XXZ-model with easy-axis anisotropy is predicted to have a peculiar ground state that can be viewed as a $\sqrt{3}\times\sqrt{3}$ “spin-supersolid” [3, 4, 5], and a series of quantum phases in applied fields. The latter include a colinear “up-up-down” ($uud$) “spin-solid” state corresponding that is an $m=1/3$ magnetization plateau. Not much is known about excitations in that model, even as significant progress has recently been made in understanding its easy plane counterpart. There, despite the presence of long range order, the excitations are nothing like those predicted by semiclassical spin wave theory (SWT) [6, 7, 8, 9]. Instead, they are dominated by bound states and continua of partially-free fractional excitations known as spinons [10]. This is taken as a fingerprint of proximate quantum spin-liquid states first hypothesized by Anderson [11, 12] and later found in numerous triangular-lattice models (see, for instance, Refs. 13, 14, 15). Does the easy-axis triangular AF feature similarly exotic spin dynamics?
We address this question experimentally and study the planar XXZ Ising-like antiferromagnet K2Co(SeO3)2 [16]. We show that in zero field and low temperatures the system has two-dimensional magnetic order consistent with a spin-supersolid. The low-energy spin excitation spectrum, however, is entirely dominated by a broad gapless spinon-like continuum, rather than by sharp spin wave modes. Applying a very modest external magnetic field induces a quantum critical point. Beyond that the material enters an $m=1/3$ $uud$ plateau. The spectrum is drastically reconstructed. It consists of sharp gapped excitations that are perfectly reproduced by SWT.
Figure 1: (a) Crystal structure of K2Co(SeO3)2. (b) The “Y” spin-supersolid phase predicted to be the ground state of an easy-axis XXZ triangular-lattice antiferromagnet. The cyan cones symbolize rotational degeneracy. (c) A magnetic field applied along the $c$ axis stabilized a colinear “$uud$” spin-solid phase.
[…]
Our first finding is that the magnetic state above 1 T is not a fully polarized one, but an $m=1/3$ magnetization plateau. This is obvious from the magnetization curves measured at $T=1.3$ K on a 5.4 mg single-crystal sample by the coaxial pick-up coil pulsed-field magnetometer at HLD-EMFL [19], and plotted in Fig. 2a. As in all experiments reported below the field is applied along the crystallographic $c$-axis (magnetic easy axis). Note that full saturation is reached only at $\mu_{0}H_{\text{sat}}\approx 22$ T. The difference between the up-sweep and down-sweep curves is instrumental and due to magnetocaloric effect [20]. The same plot shows Faraday balance magnetometry data [21] collected at $190$ mK. The low-temperature value for the start of the plateau phase is $\mu_{0}H_{c}\approx 0.8(1)$ T. Assuming the nearest-neighbor XXZ model, we can link $H_{\text{sat}}$ to the exchange constants for the in- and out-of-plane spin components: $g\mu_{B}\mu_{0}H_{\text{sat}}=3S(J_{xy}+2J_{zz})$, an exact SWT result. From the theoretical ground state phase diagram of the triangular-lattice XXZ model [2, 4], we can thereby estimate $J_{zz}\approx 3.1(3)$ meV and $\alpha=J_{xy}/J_{zz}=0.08(1)$.
A more detailed phase diagram of K2Co(SeO3)2 was established in calorimetric experiments on a Quantum Design PPMS with a dilution refrigerator insert using a 0.05 mg single-crystal sample. The result is presented in a $C/T$ specific heat plot in Fig. 2b [20]. The sharp high-temperature lambda-anomaly reported in [16] is also visible in our data (dotted line). This phase boundary has a well-defined endpoint at ($\mu_{0}H_{0}\approx 1.05$ T, $T_{0}\approx 4.5$ K). At lower field this anomaly is absent. Instead there is a broad crossover line (dashed) that extends down to zero field and corresponds to the bump at $T\sim 1$ K. An additional phase boundary marked by somewhat rounded specific heat maxima is detected at low temperatures. It has the shape of a “dome” of some ordered phase with a zero-field ordering temperature $T_{c}\approx 0.35$ K, and terminates in a quantum critical point $\mu_{0}H_{\text{c}}=0.8$ T, i.e., right at the observed boundary of the magnetization plateau. Overall, the phase diagram is qualitatively similar to that theoretically predicted for Na2BaCo(PO4)2 [5].
Figure 2: (a) Magnetization measured in K2Co(SeO3)2 in a pulsed magnetic field along the $c$ axis at $T=1.3$ K. Symbols: Faraday balance magnetometry at $T=190$ mK. Dashed lines indicate full and $1/3$ pseudospin polarization. (b) False-color plot of specific heat $C/T$ measured as a function of temperature and field. The sharp high-temperature lambda anomaly (dotted line) discussed in [16] has an endpoint (diamond). The dashed line is a broad crossover. The solid line encloses a phase pocket with an ordering temperature $T_{c}$. At $T\rightarrow 0$ there is a quantum critical point at $H_{c}$. Circles indicate the positions of local specific heat maxima in temperature scans measured at a constant field.
Next we performed neutron diffraction experiments at the DMC powder instrument and ZEBRA lifting-counter diffractometer at Paul Scherrer Insitut, as well as the AMATERAS time-of-flight (TOF) spectrometer at J-PARC [22] . In the latter we utilized 75 and 425 mg single-crystal samples, respectively. Sample environment was a 3He-4He dilution refrigerator on AMATERAS and a standard 4He cryostat on ZEBRA. In both cases a cryomagnet was used to produce a magnetic field along the $c$ axis. Elastic scattering intensity measured at $T=70$ mK in zero field in the $(h,k,0)$ plane using the TOF technique with $E_{\text{i}}=7.73$ meV incident-energy neutrons is shown in Fig. 3a. A practically identical diffraction pattern is seen in a $\mu_{0}H=1.5$ T field [20]. Both demonstrate magnetic Bragg scattering indexed by a $(1/3,1/3)$ 2D propagation vector (arrows). This corresponds to a $\sqrt{3}\times\sqrt{3}$ magnetic unit cell. In zero field this scattering sets in without any apparent phase transition (Fig. 3c, ZEBRA data), and simply increases steadily upon cooling below about 10 K. At higher fields it follows a more typical order-parameter-like curve with a transition exactly corresponding to the specific heat lambda-anomaly [20].
The observed magnetic scattering is resolution-limited in the $(h,k)$ plane. Fig. 3b show $l$-scans averaged over two sets of equivalent Bragg reflections [20]. The very broad peaks indicate only short-range correlations along $c$, of approximately AF character. The scans were fit to pairs of Voigt functions centered at $l=\pm 1/2$ (solid lines in Fig. 3b), the Gaussian component fixed at the experimental resolution [20]. From the Lorentzian width we extract the inter-plane correlation length $\lambda\approx 11(1)$, or just 0.6 lattice units. At $\mu_{0}H=1.5$ T the correlation length is only slightly larger [20]. A more detailed analysis is hardly possible given the narrow $l$-coverage.
Figure 3: Elastic scattering measured in K2Co(SeO3)2 in zero field. (a) TOF data at $T=70$ mK in the $(h,k,0)$ plane reveal $(1/3,1/3)$-type two-dimensional order (arrows). Circles are structural Bragg peaks. Solid lines are 2D Brillouin zone boundaries for a single triangular plane. “A” and “B” are two sets of equivalent Bragg points. (b) Background-subtracterd $l$-scans averaged over each set. Solid lines are fits as described in the text. The shaded Gaussian represents experimental resolution as measured on the (-2,1,0) Bragg peak. (c) Temperature dependence of diffraction intensity at $(1/3,1/3,0)$ measured on ZEBRA. The dashed line is the background level measured just off the $(1/3,1/3)$ position.
Only two sets of equivalent Bragg positions are reachable outside of the strong background region at low-$q$. Since the $uud$ and supersolid phases share the same propagation vector, a determination of the 2D magnetic structures does not appear feasible. To get a rough idea of the ordered moment in zero field, we compared the fitted Lorentzian intensities to those determined in $\mu_{0}H=1.5$ T and assumed the latter to be a $uud$ colinear structure. This procedure yields the estimate 2.6(3) $\mu_{\text{B}}$ per Co2+, substantially reduced compared to 3.95 $\mu_{\text{B}}$ at saturation (Fig. 2a).
We can seek further guidance from theory. The ground state of the classical XXZ easy-axis triangular antiferromagnet is a 3-sublattice planar “Y” structure [2], as shown in Fig. 1b. All spins on one sublattice are pointing “down”. On the two other sublattices they are canted away from the “up” direction forming an angle $\phi<2\pi/3$ between them. This is a spin density wave for the $z$-components with a $\sqrt{3}\times\sqrt{3}$ unit cell, the “down” spins forming a honeycomb. It thereby breaks a $\mathbb{Z}_{3}$ discrete translation symmetry. Crucially, the in-plane spin components of the Y-structure also breaks continuous $SO(2)$ rotational symmetry around the $z$ axis, which allows for gapless Goldstone modes.
In applied fields theory predicts a transition to the colinear $uud$ plateau phase illustrated in Fig. 1c. Here rotational symmetry around the $z$ axis is preserved and the magnetic excitation spectrum is gapped. Upon decreasing the field, the transition to the Y-phase occurs when the lowest gapped magnon softens at some critical field $H_{c}$. As realized by Batyev and Braginski sense [23] this transition is mapped on a Bose-Einstein condensation of magnons. Since it occurs in a “spin-solid” state with spontaneously broken translation symmetry, the low-field state is a “spin-supersolid” in the Andreev sense [24, 25].
For quantum spins the situation is qualitatively similar [3, 4]. This prompts us to hypothesize the following picture, that in many ways echoes the behavior predicted theoretically for Na2BaCo(PO4)2 [5]. The robust $(1/3,1/3)$ 2D magnetic correlations observed experimentally in K2Co(SeO3)2 are primarily due to the $z$ components of spins, which are similarly ordered in both $uud$- and Y-phases. They form already at $T\gg T_{c}$ and represent the spin solid. In zero field $T_{c}$ then corresponds to long-range or BKT ordering of the transverse spin components, resulting in a compressible spin-supersolid with a gapless spectrum. A BEC-type solid-supersolid quantum phase transition occurs at $H_{c}$. The main differences with the theory on Na2BaCo(PO4)2 are the presence of the end-point at $(H_{0},T_{0})$ in K2Co(SeO3)2, a much wider plateau, and a much higher saturation field.
Figure 4: False-color plots of inelastic neutron intensity measured in K2Co(SeO3)2 at $T=70$ mK in zero field (a and c), and in $\mu_{0}H=1.5$ T (b, note the ×$\times 6$ scale factor). No background is subtracted. The low-energy scattering below 0.3 meV in (b) is spurious, originating from the sample environment. It is also present in (a) but not visible due to a different color range. Panels (d,e) as well as white lines in (a,b) show simulations based on linear SWT. The red line in (d) is the lower bound of the 2-magnon continuum. (f) Equal-time structure factor obtained by integrating the intensity in the range 0.1-0.9 meV.
The emerging picture is consistent with measurements of the magnetic excitations. In Fig. 4 we show false-color neutron intensity plots of energy-momentum cuts (a,b) and energy-integrated cuts (c,f) through the data collected at the AMATERAS spectrometer using $E_{i}=2.24$ meV neutrons. More cuts are shown in the Supplement, where the integration procedure is also explained. The data in (a, c and f) are measured in zero field. Fig. 4b shows that measured at $\mu_{0}H=1.5$ T. The difference between zero field and 1.5 T data could hardly be more vivid. In the former case excitations are gapless. The spectrum is a broad continuum with no coherent modes, reminiscent of multi-spinon continua in quantum spin chains [26]. In contrast, at 1.5 T the excitations are all resolution limited.
The contrast is made even more obvious if we compare the observations with linear SWT calculations. We based these on a simple nearest-neighbor XXZ model. All simulations were performed using the SPIN-W software package [27]. First, we used simplistic Gaussian fits to the data to quantify the measured dispersion relation in the $uud$ phase. The result was then fit to linear SWT to obtain the gyromagnetic ratio $g=8.34(6)$ and exchange constants $\tilde{J}_{zz}=0.68(6)$ meV and $\tilde{J}_{xy}=0.25(1)$ meV. The fit is quite unambiguous, since the exchange parameters are well defined by the bandwidths of the upper and lower branch, respectively. The dispersion computed with these values is shown as white lines in Fig. 4a,b. The corresponding intensity simulations are shown in Fig. 4d,e [20], assuming a Gaussian broadening to match experimental energy resolution.
For $\mu_{0}H=1.5$ T SWT reproduces the experimental data up to 1 meV transfer exceptionally well. It also predicts a 3rd mode at a higher energy, one is not seen in the measurement [20]. The obtained effective SWT values $\tilde{J}_{xy}$ and $\tilde{J}_{zz}$, are hugely renormalized compared to the Hamiltonian parameters $J_{xy}$ and $J_{zz}$ deduced from the magnetization curve. This emphasizes that quantum corrections to linear-SWT remain very strong. Similar renormalization has been observed in the plateau phase of triangular-lattice materials such as Ba3CoSb2O9 [6, 7, 8].
In contrast, in zero field the computed low-energy spectra are drastically different from what is actually observed, and not only in what concerns excitation widths. For spinon continua in Heisenberg spin chains [28], the computed shape of the SWT dispersion qualitatively resembles the lower bound of the continuum. Quite to the opposite, in K2Co(SeO3)2 the latter has a distinct dip at the $M$-point, similar to the “roton-like minima” in the easy-plane case [9]. In K2Co(SeO3)2, near the $M$-point, the lower bound of the continuum shows a “relativistic” dispersion with a mass gap $\Delta_{M}=0.12$ meV. This is particulary clear from the constant-energy cut shown Fig. 4c. Weak replicas of the scattering at the $K$ and $M$ points are seen throughout the zone, revealing a hidden $\sqrt{12}\times\sqrt{12}$ periodicity in the dynamic correlations. The large $M$-point intensity contributes to the equal-time structure factor (Fig. 4f) being more or less evenly distributed along the zone boundary, in agreement with numerical simulations for the XXZ model [29].
One possible explanation for the observed behavior is spin wave instability [30], which is allowed in the non-colinear “Y”- structure. This paradigm applies well to systems with rather isotropic interactions, such as Ba3MnSb2O9 [31] or Na2BaCo(PO4)2 [32]. In K2Co(SeO3)2 it is able to account for at least some observed features. In zero field the lower bound of the two-magnon continuum in SWT coincides with the lowest-energy magnon mode (red line in Fig 4d). Above that line all modes other than the lowest energy branch around the $K$ and $\Gamma$ points are prone to decay. Experimentally, this is where the scattering is particularly broad. Even near the $K$ and $\Gamma$ points, the decay of the higher-energy acoustic mode is allowed and accounts for the continuum seen above a sharp lower bound. Unfortunately, this picture provides no explanation for the energy minimum seen at the $M$-point.
A completely different view links the continuum in the spin-supersolid phase to excitations in a proximate Dirac spin liquid [33]. In this interpretation the continuum is due to multi-spinon, rather than multi-magnon states. The predicted continuum is bounded by a “relativistic” linear-dispersion cone not only near the $K$-point, but also near the $M$ and $\Gamma$ points, exactly as observed in K2Co(SeO3)2. The theoretical intensity pattern for a constant-energy cut looks remarkably similar to what we see experimentally (Fig. 4c). We can speculate that in the high-anisotropy limit the spinon picture may provide a more adequate description of the physics.
In summary, the Ising-like triangular-lattice antiferromagnet K2Co(SeO3)2, despite long-range order, demonstrates gapless continuum spin dynamics in the spin-supersolid phase. It strongly resembles that of deconfined spinons and may be indicative of a proximate Dirac spin liquid state.
Acknowledgements.
This work was Supported by a MINT grant of the Swiss National Science Foundation. AZ thanks Prof. A. Chernyshev, Dr. M. Zhitomirsky and Prof. P. Prelovsek for providing helpful feedback. Many thanks to J. Nagl for help with graphical illustrations. We acknowledge the support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL) , and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ct.qmat (EXC 2147, project ID 390858490). Data at J-PARC were collected in Experiment no. 2023B0161. While preparing this manuscript, we became aware of Ref. [34] that seems to announce independent but similar findings.